HIGH-LYING RYDBERG STATES AND IONIZATION POTENTIAL OF VINYL CHLORIDE STUDIED BY TWO-PHOTON RESONANT IONIZATION SPECTROSCOPY

JIA-LIN CHANG, JAU-CHIN SHIEH, RUNHUA LI, JEN-CHIEH WU and \underline{YIT-TSONG CHEN}, Department of Chemistry, National Taiwan University, Taipei 106, Taiwan, and Institute of Atomic and Molecular Sciences, Academia Sinica, P.O. Box 23-166, Taipei 106, Taiwan.

High-lying Rydberg states of jet-cooled vinyl chloride (C_{2}H_{3}Cl) at 7.5-10 eV have been observed using 2+1 resonanceenhanced multiphoton ionization (REMPI) spectroscopy. The vibronic transitions in the ranges of 62500-65000 cm^{−1}, 67000-70000 cm^{−1}and 70000-72500 cm^{−1}are attributed to π →3p, π →3d/nCl→3s and π →4p excitations according to the previous assignment by Williams and Cool.^{a} Comprehensive analysis for the vibronic transitions is facilitated with the calculation of Franck-Condon factors based on our recently developed theoretical method.^{b} Four Rydberg series at 9.5-10 eV are newly observed and tentatively assigned as due to the promotion of a π-electron to the ns (n = 8-13, 20), np (n = 7-11, 18), np′ (n = 6, 7, 10-18, 21) and nf (n = 6-16) Rydberg orbitals. All the four Rydberg series converge to the same limit, i.e. the ground state of vinyl chloride cation. The term values of the nf Rydberg series, fitted to Rydberg formula, provide a very accurate adiabatic ionization energy, 80718 ± 2 cm^{−1}, for vinyl chloride with a quantum defect of δ = 0.02.

aB. A. Williams and T. A. Cool, J. Phys. Chem., 97, 1270 (1993).

bA. M. Mebel, Y. -T. Chen and S. H. Lin, J. Chem. Phys., 105, 9007 (1996).